OpenStudy (anonymous):

In triangle ABC, angle B = 30°, a = 210, and b = 164. The measure of angle A to the nearest degree is ---

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@zachleestone

OpenStudy (anonymous):

You can use the law of cosines to find an answer

OpenStudy (anonymous):



OpenStudy (anonymous):

\[c ^{2}=a ^{2}+b ^{2}-2ab \cos (B)\] \[\cos^{-1}\left(\frac{ b ^{2}-a ^{2}-c ^{2} }{ -2ac }\right)=A\]

OpenStudy (anonymous):

There's probably an easier way, but this will give you your answer